# DRAN: A Distribution and Relation Adaptive Network for Spatio-temporal Forecasting

### v1

@@ Abstract
Accurate predictions of spatio-temporal systems’ states are crucial for tasks such as system management, control, and crisis prevention. However, the inherent time variance of spatio-temporal systems poses challenges to achieving accurate predictions whenever stationarity is not granted. To address non stationarity frameworks, we propose a Distribution and Relation Adaptive Network (DRAN) capable of dynamically adapting to relation and distribution changes over time. While temporal normalization and de-normalization are frequently used techniques to adapt to distribution shifts, this operation is not suitable for the spatio-temporal context as temporal normalization scales the time series of nodes and possibly disrupts the spatial relations among nodes. In order to address this problem, we develop a Spatial Factor Learner (SFL) module that enables the normalization and de-normalization process in spatio-temporal systems. To adapt to dynamic changes in spatial relationships among sensors, we propose a Dynamic-Static Fusion Learner (DSFL) module that effectively integrates features learned from both dynamic and static relations through an adaptive fusion ratio mechanism. Furthermore, we introduce a Stochastic Learner to capture the noisy components of spatio-temporal representations. Our approach outperforms state of the art methods in weather prediction and traffic flows forecasting tasks. Experimental results show that our SFL efficiently preserves spatial relationships across various temporal normalization operations. Visualizations of the learned dynamic and static relations demonstrate that DSFL can capture both local and distant relationships between nodes. Moreover, ablation studies confirm the effectiveness of each component.
@@
Spatio-temporal systems, characterized by intricate spatial interactions among nodes and varying temporal dynamics, are prevalent in various fields such as physics [1], meteorology [2, 3, 4], power grids [5, 6] and transportation [7, 8, 9]. Real world spatio-temporal systems often entail a large number of nodes, with interactions between nodes varying over time [10, 11]. The complexity of these systems makes it challenging to make efficient decisions and manage future developments based on current conditions. Therefore, there is an urgent need for more accurate spatio-temporal prediction methods to support decision-making [12].
@@
In spatio-temporal forecasting, historical time series data associated with nodes are used to predict future observations of the same nodes [13, 14]. Although numerous methods, particularly those based on deep learning have been developed [15, 16, 17, 18, 19], achieving accurate predictions remains challenging due to the time variance of the stochastic processes generating the data. This time variance can manifest in several ways, such as distribution shifts [20, 21], changes in spatial relationships among nodes [22], and changes in the nature of noise [23].
@@
Normalization and de-normalization mechanisms are commonly employed to address distribution shifts in time series forecasting [24, 25, 26, 27]. Such techniques normalize input data to achieve consistent distributions and rebuild temporal distributions during de-normalization. Approaches such as Reversible Instance Normalization (RevIN) [24] and Dish-TS [27], which apply normalization and de-normalization on instance or temporal dimensions, have demonstrated their effectiveness in adapting to distribution shifts in time series forecasting. However, these methods may not be well-suited for spatio-temporal contexts. Methods that apply temporal normalization often scale the time series of nodes using different scaling factors. As a result, spatial message propagation occurs between nodes that have been scaled differently, which does not accurately reflect real-world message propagation processes. Moreover, instance-level normalization uses the mean and standard deviation of all nodes over the input time span, addressing only coarse spatial and temporal distribution shifts, which is inadequate for adapting to the complex distributions in spatio-temporal contexts. Therefore, it is necessary to develop approaches that achieve temporal normalization while preserving spatial patterns in spatio-temporal forecasting tasks.
@@
To capture time-varying spatial relationships, some methods learn inherent static spatial relations from data [28, 29, 30], while others generate dynamic dependencies based on input windows [31, 32] . However, static relations remain fixed after training, limiting their adaptability, while dynamic relations derived from input data often exhibit instability in predictive performance. To address these limitations, some approaches jointly learn static and dynamic characteristics and fuse them with a weighted sum [33] or a fusion ratio constrained by loss functions [34], thus achieving more comprehensive spatial representation learning [35]. However, these approaches fail to account for the dynamic interplay between static and dynamic components, relying on fixed fusion ratios that result in insufficient spatio-temporal representations. Therefore, there is a pressing need for a network capable of adapting to both static and dynamic relations while fusing their features using adaptive rations.
@@
Spatio-temporal systems inherently contain uncertainties due to data inaccuracies, noise, or sudden stochastic events within the systems. Given that these uncertainties impact prediction performance, various efforts have been made to quantify them. Some studies perform interval prediction to provide confidence regions for the forecasting horizon [36, 37]. Liang et al. [38] utilize latent variables to capture the uncertainties in air quality data, enhancing the accuracy of deterministic predictions.
@@
In this paper, we propose a Distribution and Relation Adaptive Network (DRAN) that dynamically adapts to temporal variations and assesses uncertain components influenced by noise to achieve more comprehensive representation learning in spatio-temporal systems. More specifically, we develop a Spatial Factor Learner (SFL) that incorporates temporal normalization and de-normalization stages to mitigate the impact of distribution shifts while preserving spatial dependencies. To better accommodate spatio-temporal relations, we introduce a Dynamic-Static Fusion Learner (DSFL) module that fuses features learned from sample-specific dynamic and static relations using an adaptive fusion ratio mechanism. Additionally, we propose a Stochastic Learner based on a Variational Autoencoder (VAE) to estimate the noise. Experimental validations on weather and traffic systems to forecast temperature and traffic flow demonstrate that our model outperforms state-of-the-art methods. The SFL preserves the spatial distributions across various normalization methods. The learned dynamic and static relations capture non-overlapping local and distant node relations. Ablation studies further validate the effectiveness of each component. In summary, the novelty of our method is outlined as follows:
@@
We introduce a SFL module that enables a normalization and de-normalization mechanism for distribution adaptation in a spatio-temporal context. The SFL module reverses temporal normalization while ensuring that spatial distributions closely match those of the forecasting horizons.
@@
Different from previous spatio-temporal relations adaptive methods which fuse dynamic and static features with a fixed ratio [35, 39, 33, 34], we propose a DSFL module to adaptively combine features from both dynamic and static relations with a gating mechanism, enabling accommodating the changing relations of dynamic and static perspectives.
@@
We propose a framework to learn the spatio-temporal representations from the deterministic and stochastic perspective. A VAE-based Stochastic Learner is introduced to learn the noisy components of representations.
@@
The remainder of this paper is organized as follows. Section II introduces the current work on learning spatial and temporal relations and adapting distributions. Section III details our network architecture and the overall workflow. Section IV describes the experimental setups, including dataset descriptions, network configurations, training procedures, and baseline comparisons. Section V presents the numerical results for traffic flow and weather predictions, accompanied by an ablation study and visualizations of module impacts. To have a general view of the experimental results and the effectiveness of modules, we make a discussion in Section VI. Finally, we conclude the paper and outline future research directions in Section VII.
@@ II Related Works
II-A Spatial and Temporal Relation Learning
Spatio-temporal forecasting methods are designed to capture both spatial and temporal relationships from static and dynamic perspectives. Static relations are typically learned from training datasets using trainable node embeddings and adjacency matrices to depict inter-node relationships [40, 17, 41]. Conversely, other studies [42, 43, 44] construct dynamic networks by generating sample-specific relations, which can be derived either directly from each input time series based on node similarity [45, 46] or through the use of meta-learners that generate dynamic parameters [47, 48]. Additionally, some methods refine the original relationships with input features [22, 49]. Furthermore, certain approaches simultaneously learn static and dynamic relations and integrate the gathered information. For instance, Fang et al. [34] decompose spatial relations into static and dynamic components, using a min-max learning paradigm with hyper-parameters to simultaneously learn both aspects. Similarly, Li et al. [33] employ two parallel Graph Convolutional Networks (GCNs) to learn static and dynamic information, which is then fused using a weighted sum. However, these methods typically fuse the static and dynamic features at a fixed ratio.
II-B Adaptation to Distribution Shifts
@@ III-A Problem Definition
The observed variables of node at time step can be referred to as a -dimensional vector . Each observation at time step is the result of a stochastic process, drawn from a conditioned distribution , where denotes the observations from nodes at time step in the spatio-temporal framework. The probability distribution varies over time and differs across nodes, indicating that distribution shifts occur when the observations across different time periods exhibit distinct distributions, i.e.,
where is a small threshold, is a distance function that estimates the discrepancy between distributions, and and refer to two different time steps. In this work, we utilize Gaussian kernel density estimation [53] for distribution estimation and use KL divergence-as distance function.
@@ III-B Distribution Adaptation
The distribution of historical time series often diverges from that of future time series due to time variance. Figure 2 (a) demonstrates the effectiveness of temporal normalization, which aligns the historical distribution more closely with the future distribution. This alignment facilitates easier learning of the projection from the input to the forecasting horizon. Consequently, temporal normalization is both an effective and necessary operation for spatio-temporal forecasting.
While previous works [25, 27] have addressed distribution shifts in multivariate time series forecasting, these methods are not suitable for time series with spatial relations. Direct application of these stationary strategies to spatio-temporal forecasting results in degraded performance. This is because they perform normalization within the time series of each node, leading to inconsistent scaling among nodes that are spatially interconnected. Therefore, we develop a framework that learns the temporal distribution shift on each node and preserves the spatial relations of nodes, thus maintaining the efficiency of spatial layers.
To learn the deterministic part of the inputs, We firstly filter the time series by removing high-frequency noise. We then normalize the input time series of each node using the temporal mean and standard deviation calculated for each input window.
@@
In this process, , , and represent the query, key, and value of attention, respectively, each derived from the projection of the input . The activation function is applied thereafter. To reverse the spatial patterns before inputting features into the spatial layers, we aim to de-normalize the features of nodes to preserve the spatial relations. Since the temporal representation is transformed by the De-stationary attention module, the input statistics and cannot be directly applied. As illustrated in Fig. 1 (c), we employ the SFL module to generate de-normalization factors and . In detail, we learn the node-wise features of the original input time series and the temporal representation using a 1-dimensional convolution on the temporal dimension and a Linear layer. Additionally, the statistics and are processed through a linear layer for feature dimension alignment.
Subsequently, we utilize a MLP to integrate node-wise features of the input, its temporal representation, and input data statistics to generate the de-normalization factors.
Finally, spatial factors and are applied to de-normalize .
@@ III-C Dynamic-Static Fusion Learner
Static neural networks with fixed parameters are inadequate for predicting dynamic systems. We consider that the relations of spatio-temporal systems consist of both static and dynamic components. Solely relying on input time series from different periods to establish spatio-temporal relations can lead to excessive fluctuations, thereby destabilizing prediction performance. Hence, it is essential to learn both dynamic and static information simultaneously and integrate them effectively.
To capture static information for a given task, we employ a trainable task-adaptive node embedding, akin to the approach used in the Adaptive Graph Convolutional Recurrent Network (AGCRN) [40]. Accounting for temporal variations within historical windows, we employ an adaptive node embedding to encode the static features of historical time series. Dynamic features are derived from the spatial similarity of each input data through spatial attention mechanisms.
@@
where conducts concatenation operation on the last dimension of dynamic features and static features , generating the concatenated features . We then generate the gate signal by applying a Sigmoid activation function to ; FC denotes the fully connected layer. Finally, we employ Equation (12) to fuse the dynamic and static features, producing the final deterministic features , which represent the final deterministic features learned from the input time series.
@@ III-D Stochastic Learner
After obtaining the deterministic features, we utilize both backward and forward VAEs to generate the stochastic parts of the historical and forecasting time series.
Firstly, the deterministic features are fed into the latent layers to obtain the mean and standard deviation of , thereby capturing the distribution of the input data. Then, we sample the latent features from the distribution . These latent features are then processed through the reconstruction layer to map back to the stochastic components of the time series. The processes for the backward and forward Stochastic Learners are detailed below:
@@
where and are hyper-parameters that balance the importance of various loss functions, and represents the time series generated by the neural network. Both the prediction error and the reconstruction error are computed using the Mean Absolute Error (MAE).
@@
IV Experimental Setups
In this section, we detail our experimental setup, including the datasets used, the hyper-parameter settings for our networks, the baseline comparisons, and the training process.
IV-A Datasets
Weather datasets.
NYC datasets.
We use traffic flow datasets for bikes and taxis in New York, which have been preprocessed by Ji et al. [22]. These datasets are segmented into three categories: NYCBike1, NYCBike2, and NYCTaxi, recording inflows and outflows of city bikes and taxis every 30 minutes. The NYCBike1 dataset spans from April 1st, 2014, to September 30th, 2014. The NYCBike2 dataset covers the period from July 1st, 2016, to August 29th, 2016, while the NYCTaxi dataset ranges from January 1st, 2015, to March 1st, 2015. The input and output setups are consistent with those described in Ji et al. [22]. In NYCBike1, we predict the inflow and outflow of 128 grids 30 minutes ahead using historical records of 9.5 hours, with a sequence length of 19. In NYCBike2 and NYCTaxi datasets, we utilize historical time series of 17.5 hours, with a sequence length of 35. The number of grids in NYCBike2 and NYCTaxi datasets is 200.
PeMS04 and PeMS08 datasets.
The PeMS04 and PeMS08 datasets are subsets of the PeMS (PeMS Traffic Monitoring) dataset [63], which includes real-time traffic flow data collected from loop detectors on California highways. Specifically, PeMS04 contains traffic flow records from the San Francisco Bay Area, covering the period from January 1, 2018, to February 28, 2018. PeMS08 encompasses traffic data from July 1, 2016, to August 31, 2016. Both datasets are sampled at 5-minute intervals. In our forecasting task, we aim to predict traffic flow one hour ahead based on the past one-hour records. Both the input and output lengths for each prediction task are set to 12 time steps.
Application features of these datasets are shown in Table I.

### v2

@@ Abstract
Accurate predictions of spatio-temporal systems are crucial for tasks such as system management, control, and crisis prevention. However, the inherent time variance of many spatio-temporal systems poses challenges to achieving accurate predictions whenever stationarity is not granted. In order to address non-stationarity, we propose a Distribution and Relation Adaptive Network (DRAN) capable of dynamically adapting to relation and distribution changes over time. While temporal normalization and de-normalization are frequently used techniques to adapt to distribution shifts, this operation is not suitable for the spatio-temporal context as temporal normalization scales the time series of nodes and possibly disrupts the spatial relations among nodes. In order to address this problem, a Spatial Factor Learner (SFL) module is developed that enables the normalization and de-normalization process in spatio-temporal systems. To adapt to dynamic changes in spatial relationships among sensors, we propose a Dynamic-Static Fusion Learner (DSFL) module that effectively integrates features learned from both dynamic and static relations through an adaptive fusion ratio mechanism. Furthermore, we introduce a Stochastic Learner to capture the noisy components of spatio-temporal representations. Our approach outperforms state-of-the-art methods on weather prediction and traffic flow forecasting tasks. Experimental results show that our SFL efficiently preserves spatial relationships across various temporal normalization operations. Visualizations of the learned dynamic and static relations demonstrate that DSFL can capture both local and distant relationships between nodes.
@@
Spatio-temporal systems, characterized by intricate spatial interactions among senors (nodes) and rich temporal dynamics, are prevalent in various fields e. g., physics [1], meteorology [2, 3, 4], power grids [5, 6, 7] and transportation [8, 9]. Such systems often entail a large number of nodes, with interactions among nodes varying over time [10, 11]. The complexity of these systems and the time variance makes it challenging to efficiently manage future system developments and decision-making, hence requiring more accurate spatio-temporal prediction [12].
@@
In spatio-temporal forecasting, historical time series data associated with nodes are used to predict future observations [13, 14]. Although numerous effective methods, particularly those based on deep learning, have been developed [15, 16, 17], predictions remains challenging due to the time variance of the environment-system interaction. This time variance can manifest in several ways, e. g., distribution shifts [18, 19], changes in spatial relationships among nodes [20], and changes in the nature of noise [21].
@@
Normalization and de-normalization mechanisms are commonly employed to address distribution shifts in time series forecasting, particularly in scenarios that show temporal variation without strong spatial interactions [22, 23, 24, 25]. Normalization is applied to the input data to align its distribution with that of the training set, while de-normalization maps the processed results back to the original temporal distribution. However, above methods, which apply normalization and de-normalization on instance [22] or temporal dimensions [25], may not be well-suited for processing spatio-temporal data using graph neural networks that rely on relational inductive bias. In fact, temporal normalization scales the time series of nodes using different scaling factors as the temporal content evolves with time variance, hence causing erroneous scaled spatial messages propagating between nodes. Moreover, instance-level normalization uses the mean and standard deviation of all nodes over the input time span, addressing only coarse spatial and temporal distribution shifts, which is inadequate for adapting to the complex distributions in spatio-temporal contexts. Therefore, it is necessary to develop approaches that achieve temporal normalization while preserving spatial patterns in spatio-temporal forecasting tasks.
@@
Various methods have been proposed to capture time-varying spatial relationships, including learning inherent static spatial structures from data [26, 27, 28], generating dynamic dependencies based on input windows [29, 30], and jointly modeling static and dynamic components through weighted fusion [31] or loss-constrained fusion ratios [32]. However, these approaches often overlook the dynamic interplay between static and dynamic spatial components, relying on fixed fusion schemes. This can lead to suboptimal spatio-temporal representations, especially in scenarios where the relative importance of static and dynamic features varies over time. Therefore, there is a pressing need for methods capable of adapting to both static and dynamic relations.
@@
In addition, spatio-temporal systems are inherently affected by various sources of uncertainty, including data inaccuracies, sensor noise, and sudden stochastic events that might change with time. Given that these uncertainties impact prediction performance, various efforts have been made to quantify them. Some studies perform interval prediction to provide confidence regions for the forecasting horizon [33, 34]. Liang et al. [35] utilize latent variables to capture the uncertainties in air quality data, enhancing the accuracy of predictions.
@@
In this paper, we propose a Distribution and Relation Adaptive Network (DRAN) that dynamically adapts to temporal variations and assesses uncertain components influenced by noise to achieve more comprehensive representation learning in spatio-temporal systems. More specifically, we develop a Spatial Factor Learner (SFL) that incorporates temporal normalization and de-normalization stages to mitigate the impact of distribution shifts while preserving spatial dependencies. To better accommodate spatio-temporal relations, we introduce a Dynamic-Static Fusion Learner (DSFL) module that merges features learned from sample-specific dynamic and static relations using an adaptive fusion ratio mechanism. Additionally, we propose a Stochastic Learner based on a Variational Autoencoder (VAE) to estimate uncertainty. Experimental validations on weather and traffic systems demonstrate that our model outperforms state-of-the-art methods. The novelty of our method can be summarized as follows:
@@
We introduce an SFL module that enables a normalization and de-normalization mechanisms for distribution adaptation in a spatio-temporal context. The SFL module reverses temporal normalization while ensuring that spatial distributions closely match those of the forecasting horizons.
@@
Different from previous spatio-temporal relations adaptive methods which fuse dynamic and static features with a fixed ratio [36, 37, 31, 32], we propose a DSFL module to adaptively combine features from both dynamic and static relations with a gating mechanism, accommodating the changing relations.
@@
We propose a framework to learn both deterministic and stochastic spatio-temporal representations, where a VAE-based Stochastic Learner is introduced to capture the noisy components of the representations.
@@
The paper is organized as follows. Section II introduces the state of the art. Section III details our network architecture and the overall workflow. Section IV describes the experimental setups, including dataset descriptions, network configurations, training procedures, and baseline comparisons. Section V presents the numerical results for traffic flow and weather predictions, accompanied by an ablation study and visualizations of module impacts. The discussion is provided in Section VI, and the conclusion in Section VII.
@@ II Related Works
II-A Spatio-temporal forecasting
Spatio-temporal forecasting is a longstanding and important research task. Early approaches primarily employed statistical models such as Historical Average (HA) and ARIMA [38, 39, 40], which offer interpretability but struggle to capture complex spatio-temporal dependencies. With the advent of deep learning, various modules are developed to model temporal and spatial dependencies more effectively. For temporal modeling, recurrent and convolutional neural networks [41, 42, 43] show effectiveness in short-term forecasting, while transformer-based architectures and state space models [44] demonstrate strong performance in both short- and long-term forecasting. Recurrent networks update historical information through hidden states, whereas transformer-based methods use attention mechanisms to capture dependencies from time-point [45], patch [46], and frequency-domain perspectives [47, 48]. To model spatial dependencies, both explicit and implicit methods have been proposed. Explicit methods typically utilize graph-based structures to represent node-link relationships and facilitate message propagation [28]. In contrast, implicit methods encode spatial dependencies within network parameters or attention mechanisms [49, 50, 51]. Despite promising performance, many existing methods still struggle with capturing evolving spatial structures and often lack interpretability. To address these issues, decomposition-based approaches emerge as a promising direction. These methods decompose temporal signals into trend and seasonal components [52, 53], extract multi-frequency patterns [54], and learn multi-scale spatial relationships [55, 56]. In this work, we propose a framework that decomposes spatial relations into dynamic and static components while learning both deterministic and stochastic features.
II-B Static and dynamic Relation Learning
Spatio-temporal forecasting methods are designed to capture both the static and dynamic relations existing among nodes. Static relations are typically learned from training datasets using trainable embeddings and adjacency matrices to depict inter-node relationships [57, 16, 58]. Conversely, other studies [59, 60] construct dynamic networks by generating sample-specific relations, which can be derived either directly from each input time series based on node similarity [56, 61] or through the use of meta-learners that generate dynamic parameters [62, 63]. Additionally, some methods refine the original relationships with input features [20, 64]. Furthermore, certain approaches simultaneously learn static and dynamic relations and integrate the gathered information. For instance, Fang et al. [32] decompose spatial relations into static and dynamic components, using a min-max learning paradigm with hyperparameters to simultaneously learn both aspects. Similarly, Li et al. [31] employ two parallel Graph Convolutional Networks (GCNs) to learn static and dynamic information, which is then fused using a weighted sum. These methods typically fuse the static and dynamic features at a fixed ratio. In this work, we focus on the adaptive fusion of dynamic and static spatial relations.
II-C Adaptation to Distribution Shifts
@@ III-A Problem Definition
The observed variables of node at time step can be referred to as a -dimensional vector . Each observation at time step is the result of a stochastic process, drawn from a conditioned distribution , where denotes the observations from nodes at time step in the spatio-temporal framework. The probability distribution varies over time and differs across nodes, indicating that distribution shifts occur when observations across different time periods exhibit distinct distributions, i.e.,
where is a small threshold, is a distance function that estimates the discrepancy between distributions, and and refer to two different time steps. In this work, we utilize Gaussian kernel density estimation [68] for distribution estimation and use KL divergence as a distance function.
@@ III-B Distribution Adaptation
The distribution of historical time series often diverges from that of future time series due to time variance. Fig. 2 (a) demonstrates the effectiveness of temporal normalization, which aligns the historical distribution more closely with the future distribution. This alignment facilitates easier learning of the projection from the input to the forecasting horizon. Consequently, temporal normalization is both an effective and necessary operation for spatio-temporal forecasting.
While previous works [23, 25] have addressed distribution shifts in multivariate time series forecasting, these methods are not suitable for time series with spatial relations. Direct application of these stationary strategies to spatio-temporal forecasting results in degraded performance. As shown in Fig. 2 (b), the direct application of normalization (colored in green) leads to spatial distributions deviating significantly from the spatial distribution of horizon windows. This is because they perform normalization within the time series of each node, leading to inconsistent scaling among nodes that are spatially interconnected. Therefore, we develop a framework that learns the temporal distribution shift on each node and preserves the spatial relations of nodes, thus maintaining the efficiency of spatial layers.
To learn the deterministic part of the inputs, the input time series is firstly filtered by removing high-frequency noise. We then normalize the input time series of each node using the temporal mean and standard deviation calculated for each input window.
@@
In this process, , , and represent the query, key, and value of attention, respectively, each derived from the projection of the input . The activation function is applied thereafter. To reverse the spatial patterns before inputting features into the spatial layers, we aim to de-normalize the features of nodes to preserve the spatial relations. Since the temporal representation is transformed by the De-stationary attention module, the input statistics and cannot be directly applied. As illustrated in Fig. 1 (c), we employ the SFL module to generate de-normalization factors and . In detail, we learn the node-wise features of the original input time series and the temporal representation using a 1-dimensional convolution on the temporal dimension and a Linear layer. Additionally, the statistics and are processed through a linear layer for feature dimension alignment. Subsequently, we utilize a MLP to integrate node-wise features of the input, its temporal representation, and input data statistics to generate the de-normalization factors. Finally, spatial factors and are applied to de-normalize .
@@ III-C Dynamic-Static Fusion Learner
In this section, we propose the DSFL module, which simultaneously captures dynamic and static spatial information and integrates them through an adaptive fusion mechanism. To capture static information for a given task, we employ a trainable task-adaptive node embedding, akin to the approach used in the Adaptive Graph Convolutional Recurrent Network (AGCRN) [57]. Accounting for temporal variations within historical windows, we employ an adaptive node embedding to encode the static features of historical time series. Dynamic features are derived from the spatial similarity of each input data through spatial attention mechanisms.
@@
where conducts concatenation operation on the last dimension of dynamic features and static features , generating the concatenated features . We generate the gate signal by applying a FC layer to map from to , followed by a Sigmoid activation function; here, FC denotes the fully connected layer. Finally, we employ Equation (12) to fuse the dynamic and static features, producing the final deterministic features , which represent the final deterministic features learned from the input time series.
@@ III-D Stochastic Learner
After obtaining the deterministic features, we utilize both backward and forward VAEs to generate the stochastic parts of the historical and forecasting time series. Firstly, the deterministic features are fed into the latent layers to obtain the mean and standard deviation of , thereby capturing the distribution of the input data. Then, we sample the latent features from the distribution . These latent features are then processed through the reconstruction layer to map back to the stochastic components of the time series. The processes for the backward and forward Stochastic Learners are detailed below:
@@
where and are hyperparameters that balance the importance of various loss functions, and represents the time series generated by the neural network. Both the prediction error and the reconstruction error are computed using the Mean Absolute Error (MAE).
@@
IV Experimental Setups
In this section, we detail our experimental setup, including the datasets used, the hyperparameter settings for our networks, the baseline comparisons, and the training process.
IV-A Datasets
Weather datasets.
NYC datasets. We use traffic flow datasets for bikes and taxis in New York, which have been preprocessed by Ji et al. [20]. These datasets are segmented into three categories: NYCBike1, NYCBike2, and NYCTaxi, recording inflows and outflows of city bikes and taxis every 30 minutes. The NYCBike1 dataset spans from April 1st, 2014, to September 30th, 2014. The NYCBike2 dataset covers the period from July 1st, 2016, to August 29th, 2016, while the NYCTaxi dataset ranges from January 1st, 2015, to March 1st, 2015. The input and output setups are consistent with those described in Ji et al. [20]. In NYCBike1, we predict the inflow and outflow of 128 grids 30 minutes ahead using historical records of 9.5 hours, with a sequence length of 19. In NYCBike2 and NYCTaxi datasets, we utilize historical time series of 17.5 hours, with a sequence length of 35. The number of grids in NYCBike2 and NYCTaxi datasets is 200.
PeMS04 and PeMS08 datasets. The PeMS04 and PeMS08 datasets are subsets of the PeMS (PeMS Traffic Monitoring) dataset [78], which includes real-time traffic flow data collected from loop detectors on California highways. Specifically, PeMS04 contains traffic flow records from the San Francisco Bay Area, covering the period from January 1, 2018, to February 28, 2018. PeMS08 encompasses traffic data from July 1, 2016, to August 31, 2016. Both datasets are sampled at 5-minute intervals. In our forecasting task, we aim to predict traffic flow one hour ahead based on the past one-hour records. Both the input and output lengths for each prediction task are set to 12 time steps.
Application features of these datasets are shown in Table I.